# Barrow Cosmology and Big-Bang Nucleosynthesis

### v3

@@ I Introduction
BBN refers to the process that took place in the early universe, shortly after the Big Bang, when the universe was extremely hot and dense. BBN occurred roughly between the very first centiseconds after the Big-Bang ($\sim 0.01$ sec) and a few hundred seconds ($\sim 3$ minutes) after it. During BBN, the temperature of the universe allows for nuclear reactions to occur, leading to the formation of light elements such as $H$, $He$ and $Li$. As the universe expanded and cooled, these reactions slowed down and eventually stopped, resulting in the abundance of these light elements observed today. Hence BBN, together with cosmic microwave background radiation provide a strong evidence of the hot, dense state of the early universe. In general, the principles of BBN allows to impose strict constraints on various cosmological models. The main idea is to demand consistency between theoretical predictions, and BBN observational data. In this regard, the deviation of the freeze-out temperature in comparison to standard cosmological model can be calculated. This allows to compare the theoretical results of the modified cosmological models with the observational constraints on $\left|\frac{\delta T_{f}}{T_{f}}\right|$John ; Luciano ; Anish . A modified cosmological model can also affect the primordial abundances of aforementioned light elements in the early universe Anish . It is important to note that although the standard BBN theory successfully explains the observed amounts of elements like Helium and Hydrogen, it encounters significant issues when it comes to Lithium Boran . This leads to what is known as the cosmological Lithium problem, which is a highly debated topic in the modern cosmology. It is interesting to check whether this problem can be addressed in the framework of the modified cosmology or not. The aim is to investigate the the cosmological Lithium problem in the background of the modified cosmology, to check under which conditions and which values for the model parameter, it can be alleviated.
@@ II Barrow entropy and Modified Friedman Equations
where $\tilde{r}=a(t)r$, $x^{0}=t,x^{1}=r$, and $h_{\mu\nu}$=diag $(-1,a^{2}/(1-kr^{2}))$ represents the two dimensional metric. The open, flat, and closed universes corresponds to $k=0,1,-1$, respectively. The radius of the apparent horizon, which is a thermodynamically suitable horizon for the universe, is given by Sheyem
@@
$-\frac{2+\delta}{2\pi A_{0}}\left(\frac{4\pi}{A_{0}}\right)^{\delta/2}\frac{d\tilde{r}_{A}}{\tilde{r}_{A}^{3-\delta}}=\frac{1}{3}d\rho.$ (14)
@@ III.2 Primordial light elements ${}^{4}\textrm{He}$, D and Li in Barrow cosmology
- ${}^{2}\textit{H}$ abundance- Deuterium is produced from the process $n+p\rightarrow\,^{2}\text{H}+\gamma$ Following the same approach as above, Deuterium abundance can be determined from the numerical best fit in Adv , giving
@@ III.3 Discussion on the results
$0\lesssim\delta\lesssim 0.001$ (57)
@@
$0\lesssim\delta\lesssim 0.008$ (58)
@@ IV Relation between cosmic time and temperature
[…]
@@
[…]
@@
[…]
@@
Let us note that we have used the modified Friedmann equation which has been derived in the units where $\hbar=k_{B}=c=1$. On the other hand, the right hand side of Eq. (68) should have the dimension $[\,t\,]$. If we take into account all the constants in deriving the modified Friedmann equations, the effective gravitational constant in Eq. (18) takes the form
@@
Thus, one can easily check that the right side of Eq. (68) has dimension $[\,t\,]$. In early universe, when the constituents of the universe were photons, plus three species of neutrinos and three of antineutrinos, plus electrons and positrons, we have $\mathcal{N}=43/4$, so that Eq. (68) yields, in cgs units,
@@
×$\times{2-\delta\over 4}(0.278)^{1/(\delta-2)}$
@@ V Closing remarks
(ii) We have also explored the consequences of Barrow cosmology on the primordial light elements formation. The analysis of the light elements ${}^{4}\textit{He}$, ${}^{2}\textit{H}$,${}^{7}\textit{Li}$ shows that the range of variability of $\delta$, using the bounds (51), (54) and (56), does not overlap.

### v4

@@ I Introduction
BBN refers to the process that took place in the early universe, shortly after the Big Bang, when the universe was extremely hot and dense. BBN occurred roughly between the very first centiseconds after the Big-Bang ($\sim 0.01$ sec) and a few hundred seconds ($\sim 3$ minutes) after it. During BBN, the temperature of the universe allows for nuclear reactions to occur, leading to the formation of light elements such as $H$, HeHe and $Li$. As the universe expanded and cooled, these reactions slowed down and eventually stopped, resulting in the abundance of these light elements observed today. Hence BBN, together with cosmic microwave background radiation provide a strong evidence of the hot, dense state of the early universe. In general, the principles of BBN allows to impose strict constraints on various cosmological models. The main idea is to demand consistency between theoretical predictions, and BBN observational data. In this regard, the deviation of the freeze-out temperature in comparison to standard cosmological model can be calculated. This allows to compare the theoretical results of the modified cosmological models with the observational constraints on $\left|\frac{\delta T_{f}}{T_{f}}\right|$John ; Luciano ; Anish . A modified cosmological model can also affect the primordial abundances of aforementioned light elements in the early universe Anish . It is important to note that although the standard BBN theory successfully explains the observed amounts of elements like Helium and Hydrogen, it encounters significant issues when it comes to Lithium Boran . This leads to what is known as the cosmological Lithium problem, which is a highly debated topic in the modern cosmology. It is interesting to check whether this problem can be addressed in the framework of the modified cosmology or not. The aim is to investigate the the cosmological Lithium problem in the background of the modified cosmology, to check under which conditions and which values for the model parameter, it can be alleviated.
@@ II Barrow entropy and Modified Friedman Equations
where $\tilde{r}=a(t)r$, $x^{0}=t,x^{1}=r$, and $h_{\mu\nu}$=diag $(-1,a^{2}/(1-kr^{2}))$ represents the two dimensional metric. The open, flat, and closed universes corresponds to k=0,1,1k=0,1,-1, respectively. The radius of the apparent horizon, which is a thermodynamically suitable horizon for the universe, is given by Sheyem
@@
$\frac{2+\delta}{2\pi A_{0}}\left(\frac{4\pi}{A_{0}}\right)^{\delta/2}\frac{d\tilde{r}_{A}}{\tilde{r}_{A}^{3-\delta}}=\frac{1}{3}d\rho.$ (14)
@@ III.2 Primordial light elements $^{4}\textrm{He}$, D and Li in Barrow cosmology
- $^{2}\textit{H}$ abundance- Deuterium is produced from the process $n+p\rightarrow\,^{2}\text{H}+\gamma$ . Following the same approach as above, Deuterium abundance can be determined from the numerical best fit in Adv , giving
@@ III.3 Discussion on the results
$0\lesssim\delta\lesssim 0.001.$ (57)
@@
$0\lesssim\delta\lesssim 0.008.$ (58)
@@ IV Relation between cosmic time and temperature
$t=-\frac{1}{3}\int\frac{{s}^{\prime}(T)}{s(T)}\left(\frac{8\pi G_{\rm eff}}{3}\rho\right)^{1/\delta-2}dT,$ (63)
@@
$t=-\left(\frac{8\pi G_{\rm eff}}{3}\right)^{1/\delta-2}\int\frac{1}{T}\left(\frac{\mathcal{N}a_{B}T^{4}}{2}\right)^{1/\delta-2}dT.$ (66)
@@
$t=-\left(\frac{8\pi G_{\rm eff}\mathcal{N}a_{B}}{6}\right)^{1/\delta-2}\int T^{\frac{6-\delta}{\delta-2}}dT.$ (67)
@@
Let us note that we have used the modified Friedmann equation which has been derived in the units where $\hbar=k_{B}=c=1$. On the other hand, the right hand side of Eq. (68) should have the dimension [t][\,t\,]. If we take into account all the constants in deriving the modified Friedmann equations, the effective gravitational constant in Eq. (18) takes the form
@@
Thus, one can easily check that the right side of Eq. (68) has dimension [t][\,t\,]. In early universe, when the constituents of the universe were photons, plus three species of neutrinos and three of antineutrinos, plus electrons and positrons, we have $\mathcal{N}=43/4$, so that Eq. (68) yields, in cgs units,
@@
$\times{2-\delta\over 4}(0.278)^{1/(\delta-2)}$
@@ V Closing remarks
(ii) We have also explored the consequences of Barrow cosmology on the primordial light elements formation. The analysis of the light elements $^{4}\textit{He}$, $^{2}\textit{H}$$^{7}\textit{Li}$ shows that the range of variability of $\delta$, using the bounds (51), (54) and (56), does not overlap.